OpenStudy (anonymous):

elimination method 8x-y=47 x+3y=34

OpenStudy (anonymous):

8x-y=47 ....(1) x+3y=34 ....(2) (2) x=-3y+34 (2) substitute to (1) 8(-3y+34) -y =47 -24y+272-y =47 -25y=-225 y =9 y substite to (2) x+3(9) =34 x=7

OpenStudy (anonymous):

i see what you did I forgot to do -3y+34